What is the maximum product of two numbers that add to -12? What numbers yield this product?
I came up with 36; -6 and -6 but I am not sure if that is correct

Answer by Alan3354(69443)

What is the maximum product of two numbers that add to -12? What numbers yield this product?
I came up with 36; -6 and -6 but I am not sure if that is correct
--------------------
It is correct.
Call one number x. The other one is -12-x
x*(-12-x) = 36
x^2 + 12x + 36 = 0
(x+6)*(x+6) = 0
x = -6